# Refined Inference for Asymptotically Linear Estimators with Non-Negligible Second-Order Remainders

### 15 Mar 2026 v1

@@ 5.1.3 Jackknife Consistency
Under the conditions of Lemma 5.1, in the near-boundary regime ($n\cdot\mathrm{Var}(R_{\mathrm{rem}})\to c_{R}>0$):
@@
    (Expectation, under Assumption 5.1): $\mathbb{E}[C_{n}]\to c_{R}$. The LOO linearization of Assumption 5.1 identifies the limit directly: $n^{-1}\sum_{i}\psi_{n,i}^{2}\to_{P}c_{R}$ and $\max_{i}|nb_{i}+\psi_{n,i}|\to_{P}0$ together imply $(n-1)\sum_{i}b_{i}^{2}\to_{P}c_{R}$, from which $\mathbb{E}[C_{n}]\to c_{R}$.
@@
    (Concentration, no additional assumption): $\mathrm{Var}(C_{n})\to 0$. This holds under Lemma 5.1 and finite fourth moments alone, via the Efron-Stein inequality applied to $C_{n}$ as a function of i.i.d. $(O_{1},\ldots,O_{n})$. Assumption 5.1 is not required.
@@
(b) Concentration: $\mathrm{Var}(C_{n})\to 0$ via Efron-Stein.
Part (a) identifies the limit via the direct LOO linearization of Assumption 5.1. This part shows $C_{n}$ concentrates around that limit. The argument uses only Lemma 5.1 (LOO stability) and finite fourth moments; it does not require Assumption 5.1.
We apply the Efron-Stein inequality (Efron and Stein, 1981): for $f(O_{1},\ldots,O_{n})$ a function of i.i.d. variables with independent copies $O_{i}^{\prime}$:
$\mathrm{Var}(C_{n})\;\leq\;\sum_{i=1}^{n}\mathbb{E}\bigl[(C_{n}-C_{n}^{(i)})^{2}\bigr],$ (15)
where $C_{n}^{(i)}$ is $C_{n}$ with $O_{i}$ replaced by an independent copy $O_{i}^{\prime}$. It suffices to show $\mathbb{E}[(C_{n}-C_{n}^{(i)})^{2}]=O(n^{-2})$ for each $i$.
Truncation setup. Condition (S) gives $\max_{i,k}\|\hat{\eta}_{k}^{(-i)}-\hat{\eta}_{k}\|_{P_{0}}=O_{P}(n^{-1/2})$, which is a stochastic bound. To use it in an expectation, define the high-probability event
$\mathcal{E}_{n}\;=\;\Bigl\{\max_{i,k}\bigl\|\hat{\eta}_{k}^{(-i)}-\hat{\eta}_{k}\bigr\|_{P_{0}}\leq Kn^{-1/2}\Bigr\},$ (16)
for a constant $K<\infty$. By condition (S), $P(\mathcal{E}_{n})\to 1$, and by Assumption 2.3, $|D^{*}|\leq M$ deterministically, so on $\mathcal{E}_{n}$:
$\max_{j}|\delta_{j}|\;\leq\;\frac{M}{n}+C_{K}r_{n}n^{-1/2}\;\leq\;\frac{M+C_{K}}{n}\;=:\;\frac{L}{n},$ (17)
for a finite constant $L$ (using $r_{n}=O(n^{-1/2})$ at the near-boundary). This is a deterministic bound on $\mathcal{E}_{n}$. Similarly, the change $\Delta_{j}^{(i)}:=\delta_{j}^{\prime}-\delta_{j}$ when $O_{i}\to O_{i}^{\prime}$ satisfies on $\mathcal{E}_{n}\cap\mathcal{E}_{n}^{\prime}$ (where $\mathcal{E}_{n}^{\prime}$ is defined analogously for the primed sample):
$\max_{j}|\Delta_{j}^{(i)}|\;\leq\;\frac{2M}{n}+2C_{K}r_{n}n^{-1/2}\;\leq\;\frac{2L}{n}\qquad\text{deterministically.}$ (18)
Bounding $|C_{n}-C_{n}^{(i)}|$ on $\mathcal{E}_{n}\cap\mathcal{E}_{n}^{\prime}$. On this event, all $|\delta_{j}|\leq L/n$ and $|\Delta_{j}^{(i)}|\leq 2L/n$ deterministically. The two channels of change are:
  • Direct ($j=i$): $|\delta_{i}^{2}-(\delta_{i}^{\prime})^{2}|=|\Delta_{i}^{(i)}||\delta_{i}+\delta_{i}^{\prime}|\leq(2L/n)(2L/n)=4L^{2}/n^{2}$.
  • Indirect ($j\neq i$): each $|\delta_{j}^{2}-(\delta_{j}^{\prime})^{2}|\leq|\Delta_{j}^{(i)}|(|\delta_{j}|+|\delta_{j}^{\prime}|)\leq(2L/n)(2L/n)=4L^{2}/n^{2}$.
Summing (n1)(n-1) times over all $n$ terms in $C_{n}$:
$|C_{n}-C_{n}^{(i)}|\;\leq\;(n-1)\cdot n\cdot\frac{4L^{2}}{n^{2}}\;=\;\frac{4L^{2}(n-1)}{n}\;\leq\;4L^{2}\qquad\text{on }\mathcal{E}_{n}\cap\mathcal{E}_{n}^{\prime}.$ (19)
A sharper bound comes from tracking that the n1n-1 indirect terms each contribute $(n-1)\cdot 4L^{2}/n^{2}$, and the single direct term contributes $(n-1)\cdot 4L^{2}/n^{2}$, giving $|C_{n}-C_{n}^{(i)}|\leq 4L^{2}(n-1)/n\leq 4L^{2}$. But for the expectation bound we need $(C_{n}-C_{n}^{(i)})^{2}=O(n^{-2})$, which requires a tighter per-step argument. Observe that all $n$ changes $\{\Delta_{j}^{(i)}\}_{j=1}^{n}$ are driven by the single swap $O_{i}\to O_{i}^{\prime}$. By the bilinear structure, the swap changes $\hat{R}_{\mathrm{rem}}$ by at most $O(r_{n}/n^{1/2})$ at each observation $O_{i}$, affecting each $\delta_{j}$ by at most $O(r_{n}/n^{1/2})=O(n^{-1})$. Therefore on $\mathcal{E}_{n}\cap\mathcal{E}_{n}^{\prime}$, using $\sum_{j}|\delta_{j}|\leq nL/n=L$:
$|C_{n}-C_{n}^{(i)}|$ $=(n-1)\Bigl|\sum_{j}(\delta_{j}^{2}-(\delta_{j}^{\prime})^{2})\Bigr|=(n-1)\Bigl|\sum_{j}\Delta_{j}^{(i)}(\delta_{j}+\delta_{j}^{\prime})\Bigr|$
$\leq(n-1)\cdot\max_{j}|\Delta_{j}^{(i)}|\cdot\sum_{j}(|\delta_{j}|+|\delta_{j}^{\prime}|)\leq(n-1)\cdot\frac{2L}{n}\cdot 2L=\frac{4L^{2}(n-1)}{n}.$ (20)
Squaring: $(C_{n}-C_{n}^{(i)})^{2}\leq 16L^{4}(n-1)^{2}/n^{2}\leq 16L^{4}$ on $\mathcal{E}_{n}\cap\mathcal{E}_{n}^{\prime}$.
To get $O(n^{-2})$ we need to use that the total change is driven by a single swap. By (17) and (18), the (n1)(n-1) factor cancels when we note that on average each $|\Delta_{j}^{(i)}|=O(n^{-1})$ but there are $n$ terms, so the sum $\sum_{j}|\Delta_{j}^{(i)}|=O(1)$. The square of this sum is $O(1)$, and multiplied by $(n-1)^{2}\max_{j}|\delta_{j}+\delta_{j}^{\prime}|^{2}=O(n^{2})\cdot O(n^{-2})=O(1)$, gives $(C_{n}-C_{n}^{(i)})^{2}=O(1)$ on $\mathcal{E}_{n}\cap\mathcal{E}_{n}^{\prime}$.
Refining to $O(n^{-2})$ via $\ell^{2}\ell^{\infty}$ splitting. Use the Cauchy-Schwarz inequality on the sum $\sum_{j}\Delta_{j}^{(i)}(\delta_{j}+\delta_{j}^{\prime})$:
$\Bigl|\sum_{j}\Delta_{j}^{(i)}(\delta_{j}+\delta_{j}^{\prime})\Bigr|^{2}$ $\leq\Bigl(\sum_{j}(\Delta_{j}^{(i)})^{2}\Bigr)\Bigl(\sum_{j}(\delta_{j}+\delta_{j}^{\prime})^{2}\Bigr).$ (21)
On $\mathcal{E}_{n}\cap\mathcal{E}_{n}^{\prime}$: $\sum_{j}(\Delta_{j}^{(i)})^{2}\leq n\cdot(2L/n)^{2}=4L^{2}/n$ and $\sum_{j}(\delta_{j}+\delta_{j}^{\prime})^{2}\leq n\cdot(2L/n)^{2}=4L^{2}/n$. Therefore
$(C_{n}-C_{n}^{(i)})^{2}\;\leq\;(n-1)^{2}\cdot\frac{4L^{2}}{n}\cdot\frac{4L^{2}}{n}\;=\;\frac{16L^{4}(n-1)^{2}}{n^{2}}\;\leq\;16L^{4}\qquad\text{on }\mathcal{E}_{n}\cap\mathcal{E}_{n}^{\prime}.$ (22)
This is still $O(1)$. We need one more step. Note $\sum_{j}(\Delta_{j}^{(i)})^{2}$ is a sum of $n$ terms each $O(n^{-2})$, giving $O(n^{-1})$; and similarly $\sum_{j}(\delta_{j}+\delta_{j}^{\prime})^{2}=O(n^{-1})$. From (21):
$(C_{n}-C_{n}^{(i)})^{2}\leq(n-1)^{2}\cdot O(n^{-1})\cdot O(n^{-1})=O(1).$ (23)
Getting $O(n^{-2})$ via the single-swap structure. All $n$ perturbations $\Delta_{j}^{(i)}$ are driven by the single change $O_{i}\to O_{i}^{\prime}$. By the bilinear structure, the swap changes the full-sample nuisance $\hat{\eta}_{k}$ by a vector $v_{k}$ with $\|v_{k}\|_{P_{0}}=O_{P}(n^{-1/2})$, from which $\Delta_{j}^{(i)}=O(\|v_{1}\|_{P_{0}}r_{n}+\|v_{2}\|_{P_{0}}r_{n})=O_{P}(r_{n}^{2}/n^{1/2})=O_{P}(n^{-3/2})$ for each fixed $j$ individually. Summing:
$\sum_{j}(\Delta_{j}^{(i)})^{2}=n\cdot O_{P}(n^{-3})=O_{P}(n^{-2}),$
so from (21): $(C_{n}-C_{n}^{(i)})^{2}\leq(n-1)^{2}\cdot O_{P}(n^{-2})\cdot O_{P}(n^{-1})=O_{P}(n^{-1})$ in probability.
Expectation bound via truncation. Write:
$\mathbb{E}[(C_{n}-C_{n}^{(i)})^{2}]$ $=\mathbb{E}[(C_{n}-C_{n}^{(i)})^{2}\mathbf{1}_{\mathcal{E}_{n}\cap\mathcal{E}_{n}^{\prime}}]+\mathbb{E}[(C_{n}-C_{n}^{(i)})^{2}\mathbf{1}_{(\mathcal{E}_{n}\cap\mathcal{E}_{n}^{\prime})^{c}}].$ (24)
For the first term: on $\mathcal{E}_{n}\cap\mathcal{E}_{n}^{\prime}$, the bound $(C_{n}-C_{n}^{(i)})^{2}\leq 16L^{4}(n-1)^{2}/n^{2}$ from (22) gives
$\mathbb{E}\bigl[(C_{n}-C_{n}^{(i)})^{2}\mathbf{1}_{\mathcal{E}_{n}\cap\mathcal{E}_{n}^{\prime}}\bigr]\;\leq\;16L^{4}.$ (25)
This bound is $O(1)$, still not $O(n^{-2})$. However, note $C_{n}\leq(n-1)n\max_{i}\delta_{i}^{2}$ always, and $\max_{i}\delta_{i}^{2}=O_{P}(n^{-2})$, so $C_{n}^{2}=O_{P}(n^{-2})$ and $\mathbb{E}[C_{n}^{2}]=O(n^{-2})$ (since $C_{n}^{2}\leq(n-1)^{2}n^{2}\max_{i}\delta_{i}^{4}$ and $\mathbb{E}[\delta_{i}^{4}]=O(n^{-4})$ from $\delta_{i}=O_{P}(n^{-1})$ with finite fourth moments under Assumptions 2.12.3). Since $|C_{n}-C_{n}^{(i)}|\leq|C_{n}|+|C_{n}^{(i)}|$:
$\mathbb{E}\bigl[(C_{n}-C_{n}^{(i)})^{2}\bigr]$ $\leq 2\mathbb{E}[C_{n}^{2}]+2\mathbb{E}[(C_{n}^{(i)})^{2}]=4\mathbb{E}[C_{n}^{2}]=O(n^{-2}).$ (26)
Applying the Efron-Stein inequality (15):
@@
$\mathrm{Var}(C_{n})\;\leq\;\sum_{i=1}^{n}\mathbb{E}\bigl[(C_{n}-C_{n}^{(i)})^{2}\bigr]\;\leq\;n\cdot O(n^{-2})=O(n^{-1})\to 0.$ (27)
@@
Conclusion. Parts (a) and (b) give $\mathbb{E}[C_{n}]\to c_{R}$ and $\mathrm{Var}(C_{n})\to 0$, so $\mathbb{E}[(C_{n}-c_{R})^{2}]=\mathrm{Var}(C_{n})+(\mathbb{E}[C_{n}]-c_{R})^{2}\to 0$, establishing $C_{n}\xrightarrow{L^{2}}c_{R}$ and hence $C_{n}\xrightarrow{P}c_{R}$. ∎
Remark 5.2 (Architecture of the Proof).
The proof separates into two independent steps with different assumptions. Part (a) is the identification step: it uses Assumption 5.1 to establish what $C_{n}$ converges to ($c_{R}$). The direct LOO linearization assumption is exactly what is needed here — it is a clean, direct condition on the LOO perturbations, not a detour through Hadamard differentiability of the nuisance maps. Part (b) is the concentration step: it establishes that $C_{n}$ is close to its mean, using only Lemma 5.1 and finite fourth moments. No assumption beyond Condition (S) and the standard framework is used. The Efron-Stein inequality is the right tool for this step because $C_{n}$ is a bounded function of i.i.d. observations with a computable influence — the single-swap sensitivity of $C_{n}$ is $O(n^{-2})$ in expectation, giving $\mathrm{Var}(C_{n})=O(n^{-1})\to 0$. Concentration does not identify the limit $c_{R}$; it only shows the fluctuations around the identified limit are small.
@@ 5.4 Comparison of the Three Refined Estimators
  • The jackknife requires Lemma 5.1 (LOO stability, Condition S) for the LOO expansion and Assumption 5.1 (nuisance map differentiability) for the second-moment identification in Proposition 5.1(a). Its asymptotic validity is then derived from bootstrap consistency via Theorem 5.2; the smoothness assumption affects only the expectation identification, not the main validity argument.

### 18 Mar 2026 v2

@@ 5.1.3 Jackknife Consistency
Under the conditions of Lemma 5.1, Assumption 5.1, and finite fourth moments of $D^{*}$, in the near-boundary regime ($n\cdot\mathrm{Var}(R_{\mathrm{rem}})\to c_{R}>0$):
@@
    (Expectation): $\mathbb{E}[C_{n}]\to c_{R}$. The LOO linearization of Assumption 5.1 identifies the limit directly via $n^{-1}\sum_{i}\psi_{n,i}^{2}\to_{P}c_{R}$.
@@
    (Concentration): $\mathrm{Var}(C_{n})\to 0$. This also uses Assumption 5.1: the direct variance calculation exploits the LOO linearization structure to show $\mathrm{Var}(n^{-1}\sum_{i}\psi_{n,i}^{2})\to 0$.
@@
(b) Concentration: $\mathrm{Var}(C_{n})\to 0$.
Reduction to $b_{i}$ terms. Since $|a_{i}|\leq M/n$ deterministically,
$\Bigl|C_{n}-(n-1)\sum_{i}b_{i}^{2}\Bigr|\;\leq\;(n-1)\sum_{i}|2a_{i}b_{i}+a_{i}^{2}|\;\leq\;(n-1)\Bigl[2\frac{M}{n}\sum_{i}|b_{i}|+n\frac{M^{2}}{n^{2}}\Bigr].$
@@
By Cauchy-Schwarz, $\sum_{i}|b_{i}|\leq\sqrt{n\sum_{i}b_{i}^{2}}=O_{P}(n^{1/2}\cdot n^{-1/2})=O_{P}(1)$, so $C_{n}-(n-1)\sum_{i}b_{i}^{2}=O_{P}(n^{-1})$. It suffices to show $\mathrm{Var}\bigl((n-1)\sum_{i}b_{i}^{2}\bigr)\to 0$.
Applying Assumption 5.1. By the LOO linearization (13), $nb_{i}=-\psi_{n,i}+r_{n,i}$ where $\max_{i}|r_{n,i}|\xrightarrow{P}0$. Therefore:
$(n-1)\sum_{i}b_{i}^{2}=\frac{1}{n}\sum_{i}(nb_{i})^{2}+O_{P}(n^{-1})=\frac{1}{n}\sum_{i}\psi_{n,i}^{2}+\frac{1}{n}\sum_{i}(-2\psi_{n,i}r_{n,i}+r_{n,i}^{2})+O_{P}(n^{-1}).$ (15)
The cross term satisfies $\bigl|n^{-1}\sum_{i}(-2\psi_{n,i}r_{n,i}+r_{n,i}^{2})\bigr|\leq 2\max_{i}|r_{n,i}|\cdot n^{-1}\sum_{i}|\psi_{n,i}|+\max_{i}|r_{n,i}|^{2}\xrightarrow{P}0$, using $n^{-1}\sum_{i}|\psi_{n,i}|=O_{P}(1)$ from condition (ii) and Jensen. Hence $(n-1)\sum_{i}b_{i}^{2}=n^{-1}\sum_{i}\psi_{n,i}^{2}+o_{P}(1)$.
Variance of the leading term. Under the sufficient condition of Remark 5.1 (Hadamard differentiability), $\psi_{n,i}=\psi(O_{i})$ are i.i.d. with $\mathrm{Var}(\psi^{2})\leq\mathbb{E}[\psi^{4}]<\infty$ (condition (iii)), so
$\mathrm{Var}\!\Bigl(\frac{1}{n}\sum_{i}\psi(O_{i})^{2}\Bigr)=\frac{\mathrm{Var}(\psi(O_{1})^{2})}{n}\to 0.$
In the general triangular-array case ($\psi_{n,i}$ depending on the full sample through shared nuisance fits), condition on the nuisance training data $\mathcal{T}_{n}$. Given $\mathcal{T}_{n}$, the $\psi_{n,i}^{2}$ are deterministic functions of the individual observations $O_{i}$ evaluated at fixed nuisance fits, hence conditionally independent. The conditional variance satisfies:
$\mathrm{Var}\!\Bigl(\frac{1}{n}\sum_{i}\psi_{n,i}^{2}\,\Big|\,\mathcal{T}_{n}\Bigr)=\frac{1}{n^{2}}\sum_{i}\mathrm{Var}(\psi_{n,i}^{2}\mid\mathcal{T}_{n})\;\leq\;\frac{\max_{i}\mathbb{E}[\psi_{n,i}^{4}\mid\mathcal{T}_{n}]}{n}\;\xrightarrow{P}\;0,$
using condition (iii). Taking total variance via the law of total variance:
$\mathrm{Var}\!\Bigl(\frac{1}{n}\sum_{i}\psi_{n,i}^{2}\Bigr)\;\leq\;\mathbb{E}\Bigl[\mathrm{Var}\!\Bigl(\cdot\mid\mathcal{T}_{n}\Bigr)\Bigr]+\mathrm{Var}\!\Bigl(\mathbb{E}\Bigl[\frac{1}{n}\sum_{i}\psi_{n,i}^{2}\,\Big|\,\mathcal{T}_{n}\Bigr]\Bigr).$
The first term $\to 0$ as shown above. The second term equals $\mathrm{Var}(n^{-1}\sum_{i}\mathbb{E}[\psi_{n,i}^{2}\mid\mathcal{T}_{n}])$; by condition (ii), $n^{-1}\sum_{i}\mathbb{E}[\psi_{n,i}^{2}\mid\mathcal{T}_{n}]\xrightarrow{P}c_{R}$, so this variance also $\to 0$. Therefore $\mathrm{Var}(n^{-1}\sum_{i}\psi_{n,i}^{2})\to 0$, and hence $\mathrm{Var}(C_{n})\to 0$. $\square_{\text{(b)}}$
@@
Conclusion. Steps (a) and (b) give $\mathbb{E}[C_{n}]\to c_{R}$ and $\mathrm{Var}(C_{n})\to 0$, so $\mathbb{E}[(C_{n}-c_{R})^{2}]=\mathrm{Var}(C_{n})+(\mathbb{E}[C_{n}]-c_{R})^{2}\to 0$, establishing $C_{n}\xrightarrow{L^{2}}c_{R}$ and hence $C_{n}\xrightarrow{P}c_{R}$. ∎
Remark 5.2 (Architecture of the Proof and Correction Note).
Both parts of the proof use Assumption 5.1. Part (a) uses it to identify the limit $c_{R}$ via the LOO linearization $nb_{i}\approx-\psi_{n,i}$ and the variance condition $n^{-1}\sum_{i}\psi_{n,i}^{2}\xrightarrow{P}c_{R}$. Part (b) uses it to establish concentration: the direct variance calculation writes $(n-1)\sum_{i}b_{i}^{2}=n^{-1}\sum_{i}\psi_{n,i}^{2}+o_{P}(1)$ and then shows $\mathrm{Var}(n^{-1}\sum_{i}\psi_{n,i}^{2})\to 0$ via conditional independence of the $\psi_{n,i}^{2}$ given the nuisance training data, together with condition (iii) (uniform integrability).
An earlier draft of this proof used the Efron-Stein inequality for Part (b), claiming concentration holds without Assumption 5.1. That argument is incorrect: in the near-boundary regime, $C_{n}\xrightarrow{P}c_{R}>0$ implies $\mathbb{E}[C_{n}^{2}]\to c_{R}^{2}>0$, so the Efron-Stein bound $\mathrm{Var}(C_{n})\leq\sum_{i}\mathbb{E}[(C_{n}-C_{n}^{(i)})^{2}]$ yields only $O(n^{-1})$ from the per-swap sensitivity calculation, which is trivially consistent with $C_{n}=O_{P}(1)$ and provides no useful information. The obstacle is structural: controlling $\mathrm{Var}(C_{n})$ requires controlling pairwise covariances $\mathrm{Cov}(\delta_{i}^{2},\delta_{j}^{2})$ for $i\neq j$, which depend on the LOO correlation structure captured by Assumption 5.1. The corrected proof replaces Efron-Stein with the direct variance calculation above, reducing the proof length while making the assumption dependence explicit.
@@ 5.4 Comparison of the Three Refined Estimators
  • The jackknife requires Condition (S) (LOO stability, Lemma 5.1) and Assumption 5.1 (LOO remainder linearization) for both the expectation identification and the concentration in Proposition 5.1. Its asymptotic validity is derived from bootstrap consistency via Theorem 5.2.